OpenStudy (anonymous):

(sin^2(x)+2sin(x)+1)/(sin(x)+1)

OpenStudy (mathstudent55):

The numerator is a quadratic in sin x. One way to make it easier to understand is to do a substitution. Let \(z = \sin x\) Then \(\sin^2 x = z^2\)

OpenStudy (mathstudent55):

Start by substituting z for \(\sin x\) and \(z^2\) for \(\sin ^2 x\). What do you get?

OpenStudy (anonymous):

\[(z^2+2z+1)/(z+1)\] =z+1

OpenStudy (mathstudent55):

Correct, so far. Now you need to substitute back \(\sin x \) for z.